# Extension principles for the Einstein Yang–Mills system

### v1

@@ 4.1. Extension theorem in the weighted $H^{1}$ space
Fix a number $b$ as in (1.4), and a constant $\gamma\in(b,2]$. Given datum $\alpha|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$ (namely, $\alpha|_{C_{u_{0}}\cap\{\underline{u}\leq\underline{u}_{\ast}\}}=0$), we assume the initial energy
$\|\alpha\|^{2}_{H^{1}_{w}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}$
is bounded for some $\delta>0$, and $|r-r_{\ast}|^{-\frac{\gamma}{2}}\alpha|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$. If $\delta$ is small enough, we might suppose
$\|\alpha\|^{2}_{H^{1}_{w}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}\leq\varepsilon^{2},$
where $\varepsilon$ is small enough (depending on $\delta$).
Case I: $b>0$. If $\delta$ is small enough, the solution $(r,m,\alpha)$ extends as a $C^{1}\times C^{1}\times H^{1}_{w}$ solution in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$. Moreover, for each $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$, the solution admits the energy bound
$\|\hat{w}\|^{2}_{H^{1\prime}_{w}(\underline{C}_{\underline{u}})}+\|\alpha\|^{2}_{H^{1}_{w}(C_{u})}\lesssim\varepsilon^{2}.$
In the strip $\mathcal{S}(u,\,\underline{u}_{\ast}+\delta)$, $u\in[u_{0},\,\underline{u}_{\ast})$, there is the pointwise estimate
$\xi^{-2-\gamma}r^{-2}\hat{w}^{2}+\xi^{-\gamma}\alpha^{2}+\xi^{-3-\gamma}|\partial_{u}w|^{2}+\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b}.$
Moreover, if $|(r-r_{\ast})^{1-\frac{\gamma}{2}}\alpha^{\prime}|_{L^{\infty}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}$ is bounded, then
$\xi^{2-\gamma}(\alpha^{\prime})^{2}\lesssim r^{b},\quad\text{in}\,\,\mathcal{S}(u,\,\underline{u}_{\ast}+\delta),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$ (4.3)
Case II: $b=0$. If $\delta$ is small enough, the solution $(r,m,\alpha)$ extends as a $C^{1}\times C^{0}\times H^{1}_{w}$ solution in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$. The estimates as in Case I, except the $L^{\infty}$ estimate for $\partial_{u}w$, hold as well with $b=0$.
Remark 4.2.
Since the data $\alpha|_{C_{u_{0}}}\equiv 0$ in $\underline{u}\leq\underline{u}_{\ast}$, the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is flat and hence
$m=0,\quad\text{on}\,\,\mathcal{D}(u_{0};\,\underline{u}_{\ast}).$
Therefore, it suffices to establish the energy estimates in the strip $\mathcal{S}(u,\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$.
In the proof, we have made the following continuity assumptions on the flat boundary cone $\underline{C}_{\underline{u}_{\ast}}$. Since the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is flat and hence
@@
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{-}}\int_{\underline{C}_{\underline{u}}}\left(|\xi|^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{r^{1+b}}+|\xi|^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)\mathrm{d}u=0,$
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{-}}|\xi|^{-1-\gamma}\frac{\mu}{r^{b}}+\xi^{-2-\gamma}\frac{\Phi^{2}}{r^{b}}+\xi^{-3-\gamma}\frac{(\partial_{u}w)^{2}}{r^{b}}=0.$
We assume that these quantities are continuous at $\underline{u}=\underline{u}_{\ast}$. That is, we assume
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{+}}\int_{\underline{C}_{\underline{u}}}\left(|\xi|^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{r^{1+b}}+|\xi|^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)\mathrm{d}u=0,$
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{+}}|\xi|^{-1-\gamma}\frac{\mu}{r^{b}}+\xi^{-2-\gamma}\frac{\Phi^{2}}{r^{b}}+\xi^{-3-\gamma}\frac{(\partial_{u}w)^{2}}{r^{b}}=0.$
Remark 4.3.
The estimate (4.3) indicates the singularity emanating from the initial sphere $S_{\underline{u}_{\ast},u_{0}}$ (with singular $\alpha^{\prime}$) is propagated along the incoming cone $\underline{C}_{\underline{u}_{\ast}}$.
We intend to carry out the proof of Theorem 4.1 in the remainder of this section.
4.1.1. Bootstrap assumptions
Let $\Lambda$ be a large constant that will be determined (depending on the initial energy). We suppose in $\mathcal{S}$ there are
$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}+\mu\leq{}$ $\varepsilon^{2}\Lambda^{2},$ (4.4a)
$|rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}|\leq{}$ $\varepsilon^{2}\Lambda^{2}.$ (4.4b)
It follows from the bootstrap assumption for $\hat{w}$ (4.4a) that
$w^{2}\lesssim 1+\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{S}.$ (4.5)
Using (2.4b) and the bootstrap assumption for $\mu$, we have, as stated in (3.9), that
$\underline{h}<0,\quad h>0,\quad\text{in}\,\,\mathcal{S}.$
In regard to the weight $\xi$, we list some basic facts followed from the bootstrap assumptions.
Proposition 4.4.
We have the identities concerning the derivatives of $\xi$
$r\partial_{\underline{u}}\xi=(1-\xi)\Omega^{2}h,$
$rh\frac{\partial_{u}\xi}{\xi}+\underline{h}h=rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}.$
As a result, under the bootstrap assumption (4.4a), there is
$0\leq\xi<1,\quad\partial_{u}\xi\geq 0,\quad\partial_{\underline{u}}\xi>0,\quad\text{in}\,\,\mathcal{S}.$
@@ 4.1.1. Bootstrap assumptions
For any $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$, there is always, in the strip $\mathcal{S}$, $r\geq r_{\ast}>0$ since $h>0$, and thus
$10\leq\xi<$ 1 .
Moreover, it follows that
$\partial_{\underline{u}}\xi=\frac{\Omega^{2}hr_{\ast}}{r^{2}}=(1-\xi)\frac{\Omega^{2}h}{r}>0.$
A computation yields $\partial_{u}\xi=\frac{\underline{h}-\underline{h}_{\ast}}{r}-\frac{\underline{h}\xi}{r}$, and hence
$\frac{\partial_{u}\xi}{\xi}=\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}-\frac{\underline{h}}{r}.$
Alternatively, we have the identity
$\partial_{u}\xi=\frac{-\underline{h}_{\ast}}{r}+\frac{\underline{h}r_{\ast}}{r^{2}}=-\frac{r_{\ast}}{r}\left(\frac{\underline{h}_{\ast}}{r_{\ast}}-\frac{\underline{h}}{r}\right).$
It is suggested from this formulation that $\partial_{u}\xi$ will be nonnegative if we can show that $\frac{\underline{h}}{r}$ is increasing in $\underline{u}$. In fact, using the equation (2.4c) and the bootstrap assumption (4.4a),
$\partial_{\underline{u}}\frac{\underline{h}}{r}=\frac{\partial_{\underline{u}}\underline{h}}{r}-\frac{\Omega^{2}h\underline{h}}{r^{2}}=-\left(\mu-\Phi^{2}\right)\frac{\Omega^{2}}{r^{2}}+(1-\mu)\frac{\Omega^{2}}{r^{2}}>0,$
and we finish the proof. ∎
4.1.2. Weighted energy estimates
To begin with, we multiply $\Omega^{-2}h^{-1}r^{-1-b}\xi^{-1-\gamma}$ on (3.12) and follow the procedures leading to (3.1.4). Consequently, we obtain an identity in the strip $\mathcal{S}$ for any $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$,
$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}-(1+b)\frac{\alpha^{2}}{r^{2+b}}+b\mu\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}\left(\partial_{u}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}+\partial_{\underline{u}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
Note that, the energy on the initial cone $\underline{C}_{\underline{u}_{\ast}}$, $\int_{\underline{C}_{\underline{u}_{\ast}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$, vanishes due to the continuity assumption on $\underline{C}_{\underline{u}_{\ast}}$, as shown in Remark 4.2.
With the help of Proposition 4.4, the above energy identity is further calculated as
$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{2+b-\gamma}{2}\frac{\Phi^{2}}{r^{2+b}}+(\gamma-b)\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(1+\gamma)\Omega^{2}\xi^{-2-\gamma}\frac{\Phi^{2}}{2r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$-\iint_{\mathcal{S}}(1+\gamma)\xi^{-1-\gamma}\Omega^{2}\left(rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\right)\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}.$
Since $\gamma\in(b,2]$, terms on the left hand side are all positive. At the same time, due to the bootstrap assumptions (4.4a)–(4.4b), the spacetime integrals on the right hand side have extra smallness and hence can be absorbed by the positive term $\iint_{\mathcal{S}}(\gamma-b)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left side. Therefore, we conclude
$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{2+b-\gamma}{2}\frac{\Phi^{2}}{r^{2+b}}+\frac{\gamma-b}{2}\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(1+\gamma)\xi^{-2-\gamma}\Omega^{2}\frac{\Phi^{2}}{2r^{2+b}}\lesssim\int_{C_{u_{0}}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}.$ (4.6)
Meanwhile, multiplying $\xi^{-3-\gamma}hr^{-1-b}$ on (3.14) and following the proof leading to (3.1.4), we derive another energy identity
$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}-\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left(2\alpha^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\left(2+\frac{b}{2}\right)\mu\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}\left(\partial_{\underline{u}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}+\partial_{u}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
Note that, we have used the fact that the energy $\int_{\underline{C}_{\underline{u}_{\ast}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u$ vanishes, in view of Remark 4.2. Then substituting the calculations in Proposition 4.4 into the last line of the above identity, we have
$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\frac{\gamma-b}{2}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-4-\gamma}\left(1-\xi\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}2\xi^{-3-\gamma}\Omega^{2}\alpha^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-3-\gamma}\Omega^{2}\left(\mu+\Phi^{2}+\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}\right)\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}.$
By the bootstrap assumptions (4.4a)–(4.4b), the spacetime integrals on the right hand side are in fact small error terms and can be absorbed by $\iint_{\mathcal{D}}\frac{\gamma-b}{2}\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left side. Consequently, we achieve the following energy bound
$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\frac{\gamma-b}{4}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-4-\gamma}\left(1-\xi\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
In particular, the energy estimates (4.1.2) and (4.1.2) indicates that
$\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}+\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}\lesssim\|\alpha\|^{2}_{H^{0}_{w}(C_{u_{0}})}\leq\varepsilon^{2}.$
This completes the first order weighted energy estimate.
The second order weighted energy estimate follows from multiplying $2\xi^{1-\gamma}r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}$ on (3.1.4), and modifying the proof leading to (3.1.4). Moreover, using Proposition 4.4 and the bootstrap assumptions (4.4a)–(4.4b), we obtain

### v2

@@ 4. Local existence theorem
Let $\underline{u}_{1}\in(u_{0},\,+\infty)$ be a finite constant. Suppose on the initial slice $C_{u_{0}}$, the energy is bounded
$\int_{C_{u_{0}}}r^{-1}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r<I_{1}.$
i) There exists $u_{\ast}>u_{0}$ such that the SSEYM system admits a unique solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$ in the domain $\mathcal{D}(u_{0},u_{\ast};\underline{u}_{1})$, satisfying the bound,
$\int_{C_{u}}r^{-1}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\lesssim I_{1},$
@@
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+h^{2}(\partial_{u}\hat{w})^{2}\right)\mathrm{d}u\lesssim I_{1},$
$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}+\mu\lesssim I_{1},$
ii) The solution depends continuously on the initial data in the following sense. Suppose $\left(r_{n},\,m_{n},\,\hat{w}_{n}\right)$ and $\left(r_{m},\,m_{m},\,\hat{w}_{m}\right)$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively, then
$\int_{C_{u}}\left(\frac{|\hat{w}_{n}-\hat{w}_{m}|^{2}}{r^{3}_{n}}+\frac{|\alpha_{n}-\alpha_{m}|^{2}}{r_{n}}+\frac{|\alpha^{\prime}_{n}-\alpha^{\prime}_{m}|^{2}}{r_{n}}\right)\mathrm{d}r_{n}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h_{n}^{-1}}{r_{n}}\left(\frac{\left(\hat{w}_{n}-\hat{w}_{m}\right)^{2}}{r_{n}^{2}}+\left(\alpha_{n}-\alpha_{m}\right)^{2}+h_{n}^{2}(\partial_{u}\hat{w}_{n}-\partial_{u}\hat{w}_{m})^{2}\right)\mathrm{d}u$
$\lesssim{}$ $\int_{C_{u_{0}}}\left(\frac{|\hat{w}_{0n}-\hat{w}_{0m}|^{2}}{r^{3}_{n}}+\frac{|\alpha_{0n}-\alpha_{0m}|^{2}}{r_{n}}+\frac{|\alpha^{\prime}_{0n}-\alpha^{\prime}_{0m}|^{2}}{r_{n}}\right)\mathrm{d}r_{n}.$
@@
It suffices to prove the local existence theorem near the axis, which essentially reduces the problem to a small-data regime. Away from the axis, local existence follows in a standard way.
Choose $\underline{u}_{1}>u_{0}$ such that $\underline{u}_{1}-u_{0}$ is sufficiently small, ensuring that the initial energy remains sufficiently small. We denote the size of the initial data by $\varepsilon^{2}I_{1}$, where $\varepsilon^{2}$ is small enough, namely,